The function f: R^2 to R , defined by f(x,y)= 1-y^2+x^2, has an extremum at (0,0).

True or false with full explanation


1
Expert's answer
2021-07-06T13:52:54-0400

ANSWER.

The function "f(x,y)=1-{ y }^{ 2 }+{ x }^{ 2 }" has no an extremum at the point (0,0).

EXPLANATION.

Since "f(0,y)=1-{ y }^{ 2 }\le 1=f(0,0)" and 

"f(x,0)=1+{ x }^{ 2 }\ge 1=f(0,0)"

 then by the definition of the extremum of the function "f" has no extremum at the point (0,0).